# GPFL: A Gradient Projection-Based Client Selection Framework for Efficient Federated Learning

### v1

@@ Abstract
Federated learning client selection is crucial for determining participant clients while balancing model accuracy and communication efficiency. Existing methods have limitations in handling data heterogeneity, computational burdens, and independent client treatment. To address these challenges, we propose GPFL, which measures client value by comparing local and global descent directions. We also employ an Exploit-Explore mechanism to enhance performance. Experimental results on FEMINST and CIFAR-10 datasets demonstrate that GPFL outperforms baselines in Non-IID scenarios, achieving over 9% improvement in FEMINST test accuracy. Moreover, GPFL exhibits shorter computation times through pre-selection and parameter reuse in federated learning.
1 Introduction
@@
Existing client selection algorithms for FL have limitations in addressing challenges of data quality evaluation, resource consumption minimization, and accounting for client interdependence. In terms of data quality evaluation, previous approaches have relied on metrics such as the Shapley value and data size Nagalapatti and Narayanam (2021). However, computing the Shapley value is computationally expensive and impractical for scenarios with a large number of clients. Consequently, many algorithms have resorted to assessing data quality solely based on data size, disregarding the Non-IID nature of client data. This approach inaccurately measures data quality by considering quantity alone. In the context of resource consumption, many existing client selection algorithms employ a post-selection approach, which involves distributing the model to each client for local computations of gradients and other relevant information Balakrishnan et al. (2022). However, this method incurs significant communication and computation costs. The parameter server must distribute the global model to each client, and each client carries an additional computational burden to enhance client selection accuracy. Regarding the third issue, lots of previous studies have overlooked the interdependence among clients in FL, treating each client as an independent contributor and aggregating their gradients Le et al. (2021). However, the presence of dependencies or correlations between different clients can have a significant impact on the model accuracy.
@@
2 Federated Learning with Client Selection
@@
    Client selection: There are two main approaches to client selection in FL, namely pre-selection methods like our designed method and FedCor Tang et al. (2022), and post-selection methods like DivFL Balakrishnan et al. (2022). In the pre-selection mechanism, the server first identifies suitable clients based on information from the previous round and then distributes the updated global model to those selected clients. Each of these clients computes the gradient using their local data and uploads it to the server. On the other hand, in the post-selection mechanism, the server initially distributes the parameters of the global model to all clients in round t. All clients then compute gradients based on their local data and upload them to the server. The server selects suitable candidates to participate in FL training based on the submitted gradients.
@@
    Model aggregation: The server performs aggregation of the information submitted by clients and calculates the updated global model parameters. One commonly used model aggregation strategy is FedAvg Li et al. (2020). In FedAvg, during round t, client i calculates a new local model $w_{i}^{t}$ using the global model $w^{t-1}$ and its local data. Subsequently, the weight of the global model is updated as follows:
@@ 3.1 Gradient Projection
In client selection, our goal is to choose $K$ clients that have made the most significant contribution to FL. To achieve this, we have developed an indicator called gradient projection (GP). GP is based on the assumption that the closer a client’s locally calculated gradient is to the direction pointing towards the minimum loss, the greater its contribution to FL.
@@
3.2 Gradient Projection Confidence Bound
@@
3.3 Federated Learning with Gradient Projection Confidence Bound-Based Client Selection
@@
3.4 Theoretical Analysis
Assumption 1.
@@ 3.4 Theoretical Analysis
where $t$ is the number of rounds, $\tau=\frac{2\ln n}{n_{i}}$ is the exploration term, with $n_{i}$ denoting the number of times arm $i$ has been selected until round $t$.
Due to space limitations, we present a proof outline for Theorem 1 as below. A detailed proof is available in the appendix section.
@@
4 Experiment
@@
4.1 Evaluation Setup
1) Datasets: We have conducted experiments on two real-world datasets, namely CIFAR-10 Krizhevsky et al. (2009) and Federated-MNIST (FEMNIST) Caldas et al. (2018). Following Tang et al. (2022), we used three different heterogeneous data partitions: 2 shards per client (2SPC), 1 shard per client (1SPC), and Dirichlet Distribution (Dir). In the 2SPC setting, the data is sorted based on their labels and divided into 200 shards such that all data within a single shard share the same label. Similarly, the 1SPC setting is identical to the 2SPC setting, except that each client only has one shard, meaning they possess data from only one label. Finally, the Dirichlet distribution is an unbalanced data partitioning method. For more details on our experimental setup, please refer to the Appendix.
@@ 4.1 Evaluation Setup
3) Baselines: To validate our proposed GPFL, The baselines we employed for comparison encompassed the random method, as well as the latest and representative research findings, namely FedCorTang et al. (2022) and Pow-d Cho et al. (2022). FedCor first models the loss correlations between clients using a Gaussian Process, then derives a set of $K$ clients that significantly reduce the expected global loss. Pow-d demonstrates that prioritizing clients with higher local loss leads to quicker convergence of errors. The random method randomly selects $K$ clients to participate in FL training. In particular, for the client selection process, we utilized the open-source code from https://github.com/Yoruko-Tang/FedCor to obtain the client results of FedCor, Pow-d, and Random. Subsequently, we aggregated the updated models from the selected clients using FedAvg Li et al. (2020).
4.2 Result Comparison
@@
Table 1: Test Accuracy Comparison. The FEMINST and CIFAR-10 datasets were trained for a total of 500 and 2000 rounds, respectively. The final results reported at 100% represent the average test accuracy achieved during the last 10 rounds, while the values following ’$\pm$’ indicate the maximum deviation between the test accuracy of the last 10 rounds and their respective average.
(a) 1SPC
(b) 2SPC
(c) Dir
@@
In this section, we compare the performance of GPFL with existing client selection methods. To ensure a fair comparison, all of the methods utilize FedAvg for model aggregation, resulting in the primary difference in their experimental outcomes being the client selection phase. The results on FEMINIST and CIFAR-10 datasets are presented in Table 1, while Figure 3 illustrates the accuracy fluctuations of GPFL and other baseline algorithms across three distinct settings over multiple rounds.
@@
According to Table 1 and Figure 3, we have the following observations: 1) GPFL consistently surpasses other methods on the FEMINST dataset. Across various data distributions, FEMINST quickly achieves high test accuracy, and in the end, our algorithm outperforms the highest baseline by 19.89%, 9.34%, and 10.11% on 1SPC, 2SPC, and Dir distributions respectively. Additionally, GPFL exhibits the smallest fluctuation in test accuracy over the last 10 rounds, with a difference of no more than 4% between the maximum and minimum values and the mean. This demonstrates that the proposed gradient projection (GP) can accurately measure data quality in Non-IID data, and when combined with our EE mechanism, it selects an appropriate client set. 2) On the CIFAR-10 dataset, GPFL still outperforms baseline methods under 2SPC and Dir settings, but it falls short of Pow-d under the 1SPC setting. We believe this is due to the fact that classifying animals in colored images on CIFAR-10 is a more challenging task than the classification task on the MINST dataset. The highly imbalanced 1SPC data distribution makes it difficult for FedAvg to achieve good results, which is also reflected in the large fluctuations observed in all algorithms during the last 10 rounds of this setting. While Pow-d, as a post-selection method, can select based on the running results of more clients and obtain slightly better experimental results, it comes at the cost of consuming more client computing resources. 3) FedCor has demonstrated impressive performance in the early rounds of FEMINST, but its performance on CIFAR-10 is even worse than random selection. We hypothesize that this is due to FedCor’s strategy of avoiding selecting two clients with similar data. In simple tasks with sufficient data, this approach can yield good results. However, when the task becomes more complex and there is not enough data, combining this method with FedAvg makes it difficult to achieve good experimental results.
4.3 Time Comparison
(a) FEMNIST
(b) CIFAR-10

### v2

# Gradient Projection-based Client Selection for Efficient Federated Learning

@@ Abstract
Federated learning client selection involves choosing which clients in a federated learning system will participate in training a model, balancing the trade-off between model accuracy and communication efficiency. Current methods have issues such as overlooking data heterogeneity, incurring computational burden, and treating each client as an independent contributor. To address these issues, we propose Eff-FL, a pre-selection method that selects clients based on their historical contributions and incorporates exploration and exploitation mechanisms to discover optimal client combinations. Eff-FL significantly reduces computing resources and training time while achieving high training accuracy. The key idea is that we measure the value of a client to federated training by comparing its local descent direction with the global descent direction provided by the server. On the other hand, we use the exploit-explore mechanism to explore and avoid getting stuck in local optima.
I Introduction
@@
Existing client selection algorithms for FL have limitations in addressing challenges of data quality evaluation, resource consumption minimization, and accounting for client interdependence. In terms of data quality evaluation, previous approaches have relied on metrics such as the Shapley value and data size [nagalapatti2021game]. However, computing the Shapley value is computationally expensive and impractical for scenarios with a large number of clients. Consequently, many algorithms have resorted to assessing data quality solely based on data size, disregarding the Non-IID nature of client data. This approach inaccurately measures data quality by considering quantity alone. In the context of resource consumption, many existing client selection algorithms employ a post-selection approach, which involves distributing the model to each client for local computations of gradients and other relevant information [balakrishnan2022diverse]. However, this method incurs significant communication and computation costs. The parameter server must distribute the global model to each client, and each client carries an additional computational burden to enhance client selection accuracy. Regarding the third issue, lots of previous studies have overlooked the interdependence among clients in FL, treating each client as an independent contributor and aggregating their gradients [le2021incentive]. However, the presence of dependencies or correlations between different clients can have a significant impact on the model accuracy.
@@
II Federated Learning with Client Selection
@@
    Client selection: There are two main approaches to client selection in FL, namely pre-selection methods like our designed method and FedCor [tang2022fedcor], and post-selection methods like DivFL [balakrishnan2022diverse]. In the pre-selection mechanism, the server first identifies suitable clients based on information from the previous round and then distributes the updated global model to those selected clients. Each of these clients computes the gradient using their local data and uploads it to the server. On the other hand, in the post-selection mechanism, the server initially distributes the parameters of the global model to all clients in round t. All clients then compute gradients based on their local data and upload them to the server. The server selects suitable candidates to participate in FL training based on the submitted gradients.
@@
    Model aggregation: The server performs aggregation of the information submitted by clients and calculates the updated global model parameters. One commonly used model aggregation strategy is FedAvg [DBLP:conf/iclr/LiHYWZ20]. In FedAvg, during round t, client i calculates a new local model $w_{i}^{t}$ using the global model $w^{t-1}$ and its local data. Subsequently, the weight of the global model is updated as follows:
@@ III-A Gradient Projection
In client selection, our goal is to choose $K$ clients that have made the most significant contribution to FL. To achieve this, we have developed an indicator called gradient projection (GP). GP is based on the assumption that the closer a client’s locally calculated gradient is to the direction pointing towards the minimum loss, the greater its contribution to FL. We have selected Gradient Projection as our primary metric for data quality assessment over other similarity measures, based on the following reasoning: First, FL aims to optimize via gradient descent, prioritizing directional similarity over distance measures. Since determining the exact distance to the loss minimum is impractical, angle-based similarity is more relevant. Second, adhering to FL’s privacy and efficiency, Cosine Similarity can be a candidate due to its focus on vector angles. We opted for Gradient Projection over Cosine Similarity due to its comprehensive consideration of both angular direction and magnitude. Alignment of gradient descent directions suggests data relevance but training only on similar data isn’t always best. Introducing diversity helps escape local minima and achieve a global optimum. Using gradient magnitude in client selection boosts diversity, implicitly weighting client contributions, and valuing clients with significant differences from previous data, which can improve outcomes.
@@
III-B Gradient Projection Confidence Bound
@@
III-C Federated Learning with Gradient Projection Confidence Bound-Based Client Selection
@@
III-D Theoretical Analysis
{assumption}
@@ III-D Theoretical Analysis
where $t$ is the number of rounds, $\tau=\frac{2\ln n}{n_{i}}$ is the exploration term, with $n_{i}$ denoting the number of times arm $i$ has been selected until round $t$. Due to space limitations, we present a proof outline for Theorem III-D as below. A detailed proof is available in the appendix section.
@@
IV Convergence Analysis
We have the following assumptions:
{assumption}
{assumption}
The reward factors $\mu_{i}$ are bounded in the range $[0,1]$.
We aim to prove the following theorem:
{theorem}
where $t$ is the number of rounds, $\tau=\frac{2\ln n}{n_{i}}$ is the exploration term, and $n_{i}$ denotes the number of times arm $i$ has been selected until round $t$.
Proof:
In the exploration phase of the MAB algorithm, each agent $i$ is explored for $\tau=\frac{2\ln n}{n_{i}}$ rounds. We want to analyze the expected regret $\mathbb{E}[R_{t}]$ of the algorithm in this setting.
First, we consider the exploration phase and show that after $\tau$ rounds, the average value of $(\hat{\mu}_{i}^{t})^{+}$ (the upper confidence bound estimate of agent $i$’s quality) and the average value of $\hat{\mu}_{i}^{t}$ (the estimate of agent $i$’s quality) for the agents in the selected super-arm $a_{t}$ are both less than $\epsilon_{2}$ with high probability. This can be expressed as:
$\mathbb{P}\left(\frac{1}{\tau}\sum_{t=1}^{\tau}(\hat{\mu}_{i}^{t})^{+}\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{\tau}{2}\right)$
$\mathbb{P}\left(\frac{1}{\tau}\sum_{t=1}^{\tau}\hat{\mu}_{i}^{t}\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{\tau}{2}\right)$
Next, we define $\mathcal{S}_{t}$ as the set of winner nodes at round $t$. Let $\beta$ be the maximum mean quality among all agents, i.e., $\beta=\max_{i}\mu_{i}$. Let $s_{t}$ be the cardinality of the selected super-arm. If the number of agents $i$ in $\mathcal{S}_{t}$ with $(\hat{\mu}_{i}^{t})^{+}\geq\beta+\epsilon_{2}$ is at least $s_{t}/2$, we can show that the probability of the average quality of the super-arm $a_{t}$ being less than $\beta-\epsilon_{1}$ is small. This can be expressed as:
$\mathbb{P}\left(\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}(\hat{\mu}_{i}^{t})^{+}\geq\beta+\epsilon_{2}\right)\geq\frac{s_{t}}{2}$
$\mathbb{P}\left(\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}\hat{\mu}_{i}^{t}\leq\beta-\epsilon_{1}\right)\leq\exp\left(-\frac{s_{t}}{8}\right)$
Now, let’s calculate the probability $p_{t}$ of selecting the correct super-arm at round $t$ and derive a lower bound for $p_{t}$ by considering the probabilities of different events.
The event that the true best super-arm is not selected can be expressed as the event that either the average quality of the selected super-arm is less than $\beta-\epsilon_{1}$ or the average quality of a non-selected super-arm is greater than $\beta+\epsilon_{2}$. This can be written as:
$\mathbb{P}\left(\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}\hat{\mu}_{i}^{t}\leq\beta-\epsilon_{1}\text{ or }\exists i\notin\mathcal{S}_{t}:\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}(\hat{\mu}_{i}^{t})^{+}\leq\beta+\epsilon_{2}\right)$
Using the union bound, we can bound the probability of not selecting the correct super-arm as:
$\mathbb{P}\left(\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}\hat{\mu}_{i}^{t}\leq\beta-\epsilon_{1}\text{ or }\exists i\notin\mathcal{S}_{t}:\frac{1}{s_{t}}\sum_{i\in\mathcal{S}_{t}}(\hat{\mu}_{i}^{t})^{+}\leq\beta+\epsilon_{2}\right)$
$\leq\exp\left(-\frac{s_{t}}{8}\right)+\exp\left(-\frac{\tau}{2}\right)$
Let $\delta=\exp\left(-\frac{s_{t}}{8}\right)+\exp\left(-\frac{\tau}{2}\right)$. Then, the probability $p_{t}$ of selecting the correct super-arm at round $t$ is given by $p_{t}=1-\delta$.
Using the derived lower bound for $p_{t}$, we can obtain an upper bound for the expected regret $\mathbb{E}[R_{t}]$ of the MAB algorithm. Therefore, the expected regret can be bounded as:
$\mathbb{E}[R_{t}]\leq\frac{1}{p_{t}}(\beta+\epsilon_{1})\leq\frac{1}{1-\delta}(\beta+\epsilon_{1})$
Finally, we substitute the value of $\delta$ and simplify the expression to obtain the desired regret bound:
This completes the proof. ∎
V Experiment Settings
V-A Dataset and Data Distribution
We conducted federated learning experiments on two benchmark datasets: FEMNIST and CIFAR-10. Federated Extended MNIST (FEMNIST) is a dataset specifically crafted for federated learning research, offering a realistic and challenging environment for developing and testing machine learning models. It is based on the Extended MNIST dataset, which includes both handwritten digits (0-9) and letters (A-Z, a-z). The dataset is part of the LEAF benchmark suite, which includes a range of datasets and evaluation tools to facilitate reproducible research and comparison of federated learning algorithms. CIFAR-10 is a high-quality dataset for image classification tasks, consisting of 60,000 32x32 color images categorized into 10 classes, each containing 6,000 images. The original dataset is evenly split into 50,000 training images and 10,000 test images, and it includes a diverse set of objects such as airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
The partitioning of the two datasets is designed to emulate a federated learning environment, wherein each device or user contributes data originating from a distinct source. The summary statistics for these datasets are presented in Table I, which encompasses details such as the count of devices, the total number of samples, the average number of samples per device, and the associated standard deviation.
FEMINST CIFAR-10
Client Count 3,550 100
Sample Count 805,263 60,000
Sample per Client Mean 226.83 946.8
Stdev 88.94 256.04
TABLE I: Statistics of dataset
We investigate three different heterogeneous settings to analyze the effects of data partitioning and client heterogeneity on the performance of federated learning:
  1. 1.
    2 shards per client (2SPC): The data is sorted by labels and divided into 200 shards, where each shard contains samples from the same label. The shards are randomly allocated to clients, with each client receiving two shards. As all shards have the same size, the data partition is balanced, ensuring that all clients have the same dataset size. We select K = 5 clients in each round within this setting.
  2. 2.
    1 shard per client (1SPC): This setting is similar to the 2SPC setting, but each client is assigned only one shard, representing the highest level of data heterogeneity. Each client contains data from a single label, resulting in a balanced partition. In each round within this setting, we select K = 10 clients.
  3. 3.
    Dirichlet Distribution with $\zeta=0.2$ (Dir): For each client, denoted as $i$, we independently sample the data distribution $q_{i}$ from a Dirichlet distribution, which can be formulated as $q_{i}\sim\text{Dir}(\zeta p)$. Here, $p$ represents the prior label distribution, and $\zeta\in\mathbb{R}^{+}$ is the concentration parameter of the Dirichlet distribution. By combining the $q_{i}$ values for all clients, we obtain a fraction matrix $Q=[q_{1},q_{2},...,q_{n}]$. The dataset size on each client is denoted as $x=[x_{1},x_{2},...,x_{N}]^{T}$, and we determine it through a quadratic programming solution. We minimize $\lVert x\rVert_{2}$ while ensuring that the matrix product $Qx$ equals the desired distribution $d$, where $d$ represents the number of data samples for each label. This minimization helps avoid cases where the data distribution becomes overly concentrated on a small fraction of clients, as it would make the client selection problem trivial. We set the concentration parameter $\zeta$ to 0.2 to create the Dirichlet Distribution with $\zeta=0.2$ (Dir) setting, resulting in an unbalanced data partition. This setting introduces varying proportions of different labels across clients, generating heterogeneity in the dataset partition. In each round of this setting, we select five clients ($K=5$) to participate in the federated learning process.
V-B Model Construction and Hyperparameters
V-B1 Model Construction and Hyperparameters in FEMNIST
For the FEMNIST dataset, we adopt an MLP (Multi-Layer Perceptron) architecture with two hidden layers, consisting of 64 and 30 units, respectively. In all three heterogeneous settings, the local batch size is set to 64, and the number of local iterations is set to 20. We initialize the learning rate ($\eta$) as 0.005. We utilize an SGD (Stochastic Gradient Descent) optimizer with a weight decay of 0.0001 and momentum of 0.1. The data is divided among $N=100$ clients, with a participation fraction ($K$) of 10 for the 1SPC setting and 5 for the 2SPC and Dir settings.
V-B2 Model Construction and Hyperparameters in CIFAR-10
For the CIFAR-10 dataset, we employ a CNN (Convolutional Neural Network) architecture with three convolutional layers, consisting of 32, 64, and 64 kernels, respectively, followed by a fully-connected layer with 64 units. In all three heterogeneous settings, the local batch size is set to 50, and the number of local iterations is set to 40. We use a fixed learning rate ($\eta$) of 0.01 without any learning rate decay and a weight decay of 0.0003 for the SGD optimizer. The total number of clients and the client participation fraction are consistent with those used in the FEMNIST dataset.
By leveraging these specific hyperparameters and employing the respective neural models, we conducted federated learning experiments on the FEMNIST and CIFAR-10 datasets under different heterogeneous settings.
VI Experiment
@@
VI-A Evaluation Setup
1) Datasets: We have conducted experiments on two real-world datasets, namely CIFAR-10 [krizhevsky2009learning] and Federated-MNIST (FEMNIST) [caldas2018leaf]. Following [tang2022fedcor], we used three different heterogeneous data partitions: 2 shards per client (2SPC), 1 shard per client (1SPC), and Dirichlet Distribution (Dir). In the 2SPC setting, the data is sorted based on their labels and divided into 200 shards such that all data within a single shard share the same label. Similarly, the 1SPC setting is identical to the 2SPC setting, except that each client only has one shard, meaning they possess data from only one label. Finally, the Dirichlet distribution is an unbalanced data partitioning method. For more details on our experimental setup, please refer to the Appendix.
@@ VI-A Evaluation Setup
3) Baselines: To validate our proposed GPFL, The baselines we employed for comparison encompassed the random method, as well as the latest and representative research findings, namely FedCor[tang2022fedcor] and Pow-d [cho2022towards]. FedCor first models the loss correlations between clients using a Gaussian Process, then derives a set of $K$ clients that significantly reduce the expected global loss. Pow-d demonstrates that prioritizing clients with higher local loss leads to quicker convergence of errors. The random method randomly selects $K$ clients to participate in FL training. In particular, for the client selection process, we utilized the open-source code from https://github.com/Yoruko-Tang/FedCor to obtain the client results of FedCor, Pow-d, and Random. Subsequently, we aggregated the updated models from the selected clients using FedAvg [DBLP:conf/iclr/LiHYWZ20].
VI-B Result Comparison
@@
TABLE II: Test Accuracy Comparison. The FEMINST and CIFAR-10 datasets were trained for a total of 500 and 2000 rounds, respectively. The final results reported at 100% represent the average test accuracy achieved during the last 10 rounds, while the values following ’$\pm$’ indicate the maximum deviation between the test accuracy of the last 10 rounds and their respective average.
((a)) 1SPC
((b)) 2SPC
((c)) Dir
@@
In this section, we compare the performance of GPFL with existing client selection methods. To ensure a fair comparison, all of the methods utilize FedAvg for model aggregation, resulting in the primary difference in their experimental outcomes being the client selection phase. The results on FEMINIST and CIFAR-10 datasets are presented in Table II, while Figure 3 illustrates the accuracy fluctuations of GPFL and other baseline algorithms across three distinct settings over multiple rounds.
@@
According to Table II and Figure 3, we have the following observations: 1) GPFL consistently surpasses other methods on the FEMINST dataset. Across various data distributions, FEMINST quickly achieves high test accuracy, and in the end, our algorithm outperforms the highest baseline by 19.89%, 9.34%, and 10.11% on 1SPC, 2SPC, and Dir distributions respectively. Additionally, GPFL exhibits the smallest fluctuation in test accuracy over the last 10 rounds, with a difference of no more than 4% between the maximum and minimum values and the mean. This demonstrates that the proposed gradient projection (GP) can accurately measure data quality in Non-IID data, and when combined with our EE mechanism, it selects an appropriate client set. 2) On the CIFAR-10 dataset, GPFL still outperforms baseline methods under 2SPC and Dir settings, but it falls short of Pow-d under the 1SPC setting. We believe this is due to the fact that classifying animals in colored images on CIFAR-10 is a more challenging task than the classification task on the MINST dataset. The highly imbalanced 1SPC data distribution makes it difficult for FedAvg to achieve good results, which is also reflected in the large fluctuations observed in all algorithms during the last 10 rounds of this setting. While Pow-d, as a post-selection method, can select based on the running results of more clients and obtain slightly better experimental results, it comes at the cost of consuming more client computing resources. 3) FedCor has demonstrated impressive performance in the early rounds of FEMINST, but its performance on CIFAR-10 is even worse than random selection. We hypothesize that this is due to FedCor’s strategy of avoiding selecting two clients with similar data. In simple tasks with sufficient data, this approach can yield good results. However, when the task becomes more complex and there is not enough data, combining this method with FedAvg makes it difficult to achieve good experimental results.
VI-C Result Analysis
Figure 4: Number of Clients Selected At Least Once Over Rounds
((a)) GPFL
((b)) FedCor
((c)) Pow-d
((d)) Random
Figure 5: Histogram for selection frequency
In this section, we delve deeper into the analysis of our experimental results, focusing on the characteristics of client selection exhibited by various algorithms. We conducted a pair of experiments designed to evaluate these characteristics. The first experiment monitored the proportion of clients selected at least once as the number of training rounds increased. It is imperative in Federated Learning to ensure that client selection algorithms do not consistently omit certain clients, as each client should ideally be chosen early on to gauge their unique contribution effectively. This particular experiment tracks how many rounds it takes for every algorithm to include all clients in the selection process at least once. The second experiment plotted the frequency of client selections against the total number of clients, providing insight into the propensity of the algorithms towards certain clients over others.
Illustrated in Figure 4, we observe the progression of the percentage of clients selected across different training rounds. The x-axis denotes the training rounds, while the y-axis denotes the proportion of clients selected at least once per round. Our proposed GPFL algorithm impressively ensured that all clients were selected at least once within a mere 50 rounds. This performance significantly outpaced the nearly 100 rounds required by random and pow-d algorithms, with FedCor taking approximately 200 rounds to achieve equivalent coverage. The accelerated rate at which GPFL selects clients can be attributed to its strategic exploration and exploitation approach. In the early stages of training, when the element of exploitation is not yet stable, GPFL assigns a higher priority to exploration, favoring clients that have not been previously selected. This tactic enables a more rapid and comprehensive selection compared to random and pow-d methods. Such prioritization is particularly beneficial in federated learning scenarios where data heterogeneity is crucial, as it allows the model to leverage the full spectrum of available data early on, potentially reducing bias and enhancing overall performance. The prolonged selection process exhibited by FedCor—surpassing 200 rounds before including all clients—hints at a subset of non-selected clients that were overlooked during the initial rounds over an extended period.
The histogram in Figure 5 presents the results of the second experiment, illustrating the frequency of client selection versus the number of clients. Here, the x-axis represents the frequency of selections, and the y-axis denotes the quantity of clients. Across all tested baselines, a roughly normal distribution is evident with an average around 20. However, GPFL diverges with a distinctive long tail reaching up to around 500, suggesting that a small group of clients were selected considerably more frequently. This pattern indicates that GPFL strikes a balance between selecting a diverse range of clients and repeating selections for a chosen few. Conversely, ’Fedcor’ displays a long tail extending to about 120 selections, indicating a tendency to repeatedly choose a limited subset of clients. Pow-d and Random exhibit more conventional normal distributions, with Pow-d showing a flatter curve centered on (24, 7) and Random peaking around (25, 10). The distinctive feature of GPFL lies in its ability to select a broad range of clients while also favoring repetition for certain clients. This balance ensures a diverse representation in training while allowing for deeper exploration of influential client data. This strategy potentially contributes to the effectiveness of GPFL in leveraging the full spectrum of available client data for model training, leading to enhanced model robustness and generalization capability.
VI-D Time Comparison
((a)) FEMNIST
((b)) CIFAR-10